# Tales of the 2025 Los Angeles Fire: Hotwash for Public Health Concerns in Reddit via LLM-Enhanced Topic Modeling

### v2

@@ 1 Introduction
Wildfires have intensified in intensity, severity, and duration on a global scale, largely due to the effects of climate change and increased human activity Swain et al. (2025); Mansoor et al. (2022); Yue et al. (2021). These unplanned fire events result in widespread structural damage, economic losses, environmental degradation, and serious public health threats to affected communities Bowman et al. (2017); D’Evelyn et al. (2022); Rappold et al. (2017); Wang et al. (2016). The health risks associated with wildfires extend far beyond the immediate fire zones, encompassing not only physical and mental health concerns but also disparities in access to healthcare resources. Epidemiological and toxicological evidence links exposure to wildfire-derived fine particulate matter (PM2.5) to a range of symptoms, including respiratory and ocular irritation, and to adverse health outcomes such as asthma, cardiovascular diseases, negative birth outcomes, and skin inflammation Zhang et al. (2025); Lei et al. (2024). In addition, psychological impacts—such as anxiety, depression, and trauma-related disorders can result from smoke exposure, fear of loss, and evacuation-related stress Lei et al. (2024); Wettstein and Vaidyanathan (2024). These mental health effects may persist long after fires are contained and can exacerbate physical health issues, particularly among vulnerable populations including children, older adults, pregnant individuals, those with preexisting cardiopulmonary conditions, and individuals from lower socioeconomic backgrounds Lei et al. (2024); Rappold et al. (2017).
@@
When wildfires occur during irregular seasons and communities are unprepared, the consequences can be even more severe. In California’s Mediterranean climatic region, January has historically been a low-probability period for large-scale wildfires Dong et al. (2022). However, a hydroclimate whiplash Homann et al. (2022); Swain et al. (2025), driven by low humidity and hurricane-force Santa Ana winds, has induced abrupt changes from protracted wet periods to extreme dryness. This shift triggered the most devastating winter wildfires in the history. As of January 31, 2025, a total of 28 wildfires had burned 57,636 acres, resulted in 29 fatalities (pending coroner confirmation), destroyed 16,244 structures of Forestry and Protection (2025), and forced the evacuation of approximately 200,000 residents Stelloh et al. (2025). Preliminary data indicate that the Eaton and Palisades fires in Los Angeles rank second and third in structural damage, and fifth and ninth in fatalities, respectively, though these figures remain provisional California Department of Forestry and Fire Protection (2025) (CAL FIRE). Despite the Storm Prediction Center Storm Prediction Center (2025) (SPC), National Interagency Fire Center National Interagency Fire Center (2025) (NIFC), and the local National Weather Service National Weather Service (2025) (NWS) had accurately forecasted the windstorm and heightened fire risk, the actual spread and scale of wildfires far exceeded expectations.
@@
Amid rapidly evolving wildfire conditions, many individuals have turned to social media platforms—such as Twitter and Reddit—for hyperlocal information, health-related resources, and mutual aid Slavkovikj et al. (2014); Lever et al. (2022); Ma et al. (2024). While Twitter has been extensively studied in the context of disaster response, Reddit’s role in collective information sharing and community support remains comparatively underexplored Chen and Ye (2019). Traditional data sources, including clinical records and post-disaster surveys, offer valuable medical insights into wildfire-related health outcomes; however, they often suffer from limitations such as small sample sizes, limited scalability, high costs, and delayed data availability García et al. (2024). With the California fire season approaching, there is an urgent need for rapid, scalable methods to detect public health signals and uncover unmet needs in affected communities in real time.
@@
Social media platforms serve as critical sources of information during disasters, significantly supporting disaster response through the rapid dissemination of real-time updates Ye et al. (2020). Given the sheer volume of data generated on social media platforms far beyond the capabilities of manual examination, existing research leverages machine learning techniques, such as keyword-based methods or classification models, which utilize textual and metadata features, to identify and categorize relevant social media content Vongkusolkit and Huang (2021); Huang and Xiao (2015). Most recently, pretrained large language models (LLMs) have become the state-of-the-art method in natural language processing (NLP). Built on top of these pretrained embeddings, supervised and unsupervised methods are applied as post-processing strategies. Supervised techniques still require a certain quantity of annotated data for fine-tuning, which can be time-consuming in fast-moving emergencies. By contrast, unsupervised methods, such as topic modeling, can quickly identify latent topics based on semantic clustering without annotation efforts Zhou et al. (2023). Furthermore, most previous research on disaster events has considerable time lags, with findings published years after the crisis. These findings can inform long-term mitigation, but often lack value for immediate response.
@@
In order to quickly identify and extract actionable information for disaster response and management, a topic classification schema or a coding schema is needed to guide the categorization of social media datasets Vongkusolkit and Huang (2021); Verma et al. (2011); Hughes and Palen (2009). These guidance schema often include categories such as casualties, damage reports, donation requests, or public emotions. However, designing an effective classification schema and corresponding information extraction methods is complex and highly context-dependent, influenced by factors such as (1) the type of disaster event, (2) the specific goals of the study or analysis, and (3) the characteristics of the social media platform used Vongkusolkit and Huang (2021).
@@ 2 Literature Review
Topic modeling is an unsupervised NLP technique that enables the identification and organization of latent thematic structures within large text corpora. This approach is particularly valuable for managing the exponential growth of digital textual data—from emails and reports to social media posts—by converting unstructured content into structured, analyzable representations Vayansky and Kumar (2020).
@@ 2.1 Topic Modeling Algorithms
The core of many topic modeling methods is the “bag of words” paradigm, where the focus is on word frequencies rather than their order Zhang et al. (2010). This paradigm generates two critical matrices: a word-topic matrix that encapsulates topics through their associated words, and a document-topic matrix that reflects each document’s composition in terms of these topics. Such representations help extract important insights from vast corpora, especially in the absence of explicit labeling Barde and Bainwad (2017).
Historically, topic modeling has evolved significantly—from early models such as Latent Semantic Analysis (LSA) Deerwester et al. (1990) and its probabilistic extension, Probabilistic LSA (pLSA) Hofmann (1999), to more sophisticated Bayesian approaches like the Gibbs sampling algorithm for the Dirichlet multinomial mixture (GSDMM) Yin and Wang (2014) and Latent Dirichlet Allocation (LDA). LDA, for instance, leverages Dirichlet distributions to iteratively refine the assignment of words to topics based on the probability of topic occurrence in documents and word occurrence within topics Blei et al. (2003). Concurrently, Non-Negative Matrix Factorization (NMF) emerged as an alternative linear algebra-based method that decomposes high-dimensional text data into interpretable, lower-dimensional representations Lee and Seung (1999).
In recent years, the advent of transformer-based models has ushered in a new era of topic modeling with algorithms such as Top2Vec Angelov (2020) and BERTopic Grootendorst (2022). Top2Vec detects topics with assumption that each document is primarily centered on a single topic. In contrast, BERTopic performs powerful and flexible modular structure by integration of sophisticated sentence embedding techniques, dimensionality reduction, and clustering algorithms for modern text analytics.
Overall, the evolution from LSA and pLSA to GSDMM, LDA, NMF, and the transformer-based methods illustrates the dynamic development of topic modeling methodologies. Each algorithm has its own merits for applications depending on the trade-offs among scalability, interpretability, and contextual sensitivity. For example, LDA is well-suited for longer texts but is typically efficient only when applied to a relatively small number of documents. In contrast, BERTopic performs better on shorter texts but requires a large corpus to effectively capture and cluster meaningful topics Egger and Yu (2022).
@@ 2.2.1 Dataset Sources
Structured data sources, such as NASA’s Aviation Safety Reporting System (ASRS) and the ICS-209-PLUS incident database, have been used in a limited number of studies focusing on operational safety and technical failures in wildfire contexts Mbaye et al. (2023); Andrade and Walsh (2023)). While these datasets are valuable in terms of infrastructure performance and risk reporting, they are less suited for examining public sentiment, communication, and response in real time.
In contrast, most studies rely on social media data, particularly from Twitter, to analyze wildfire trends, public discourse, and crisis communication strategies. These datasets vary in size, ranging from thousands to millions of posts, and span from short-term event monitoring to multi-year longitudinal analyses Zander et al. (2023); García et al. (2024); Souza et al. (2024); Madichetty et al. (2021); Madichetty and Madisetty (2023). For instance, García et al. (2024) focused on mental health during the 2017 Tubbs Fire in California, while Zander et al. (2023) examined broader situational awareness during Australia’s “Black Summer” bushfires in 2019 and 2020.
Due to limitations associated with Twitter data—such as character constraints, low information density, credibility issues, and recent restrictions on data access—researchers have begun to explore alternative platforms. Reddit, with its longer-form user content, community-based moderation, and more accessible data policies, presents a compelling alternative for analyzing public discourse around wildfires Lever et al. (2022).
@@ 2.2.2 Topic Classification Schema
After latent topic clusters are generated, it is common to organize them into interpretable categories using domain annotations. Many studies implicitly follow two analytical rationales: situational awareness (SA) and emotional response. SA is defined as “all knowledge that is accessible and can be integrated into a coherent picture, when required, to assess and cope with a situation” Sarter and Woods (2017), derived from three-level model of perception, comprehension, and projection Endsley (1995). In the context of disaster response, SA captures real-time information such as fire progression, evacuation orders, and resource availability. While SA-based categories are prevalent across studies, shown in Table 1, they tend to vary widely ranging from three to fifteen categories depending on research goals. This inconsistency reflects a lack of unified structure for organizing real-time information in disaster response.
To capture emotional response, many studies rely on sentiment analysis Lever et al. (2022); García et al. (2024); Zander et al. (2023), which classifies corpus categorizing text into positive, negative, or neutral categories, as well as fundamental emotion such as fear, anger, or sadness. While sentiment analysis provides general emotional trends, it is often insufficient to summarize intricate expressions such as narratives of blame, resilience, or justice Ruiz-Soler (2017). Moreover, sentiment labels may be inaccurate in informal, sarcastic, or ambiguous media messages Wankhade et al. (2022).
@@
Madichetty et al. (2021) Califonia wildfires Twitter 2 weeks 15 Fine-tuned BERT
Lever et al. (2022) California Wildfire Season (2016) Twitter, Reddit 4 months 5 LDA
Ma et al. (2024) Western Wildfire Season(2020) Twitter 1 month 6 BERTopic
Zander et al. (2023) Black Summer(2019-2020) Twitter 6 months 8 GSDMM
Souza et al. (2024) Wildfires in Brazil National Park Twitter 11 years 6 BERTopic
García et al. (2024) Tubbs fire Twitter 3 weeks 8 LDA
Andrade and Walsh (2023) Wildfire incidents ICS-209-PLUS 8 years 3 BERTopic
Mbaye et al. (2023) Wildfire reports UAS 4 years 7 BERTopic
To address these limitations, we draw on the concept of Crisis Narratives from crisis communication research. Crisis Narratives (CN) refer to “the stories told about crises that carry meaning, encode lessons, and frame larger public and societal understanding of risks, warnings, and potential harm” Seeger and Sellnow (2016); Liu et al. (2020). Common narrative themes, such as blame, victimhood, heroism, and renewal, have been widely used to examine institutional rhetoric and public interpretation in disaster contexts, as shown in table 2. Despite their relevance, this framework has not yet been systematically applied to social media data in the context of wildfires. In this study, we adopt and extend CN theory to move beyond surface-level sentiment analysis, offering a more nuanced and structured lens for interpreting emotionally charged public discourse during crises.
@@
Wolfe (2016) Disaster Interview, participant observation, public meetings, media reports and organizational documents Blame and Accountability, Victim-Centered Disaster, Renewal and Growth
Liu et al. (2020) Pandemic Quantitative survey Blame, Renewal, Victim, Hero, Memorial
An (2024) Pandemic Twitter Blame, Renewal, Victim, Hero, Memorial
@@ 3.2 Data Preprocessing
We adapted a multi-step preprocessing to clean the Reddit data for analysis. In the first step, we removed special characters, HTML tags, and Markdown symbols. We excluded instances with fewer than 10 words based on the recommended effective semantic length (15-20 words) Markel and Markel (2009). Unlike traditional NLP methods that completely delete emojis and URLs, we extracted and saved 1,002 unique URLs embedded in posts and comments for rank analysis. Emojis were converted to their corresponding textual descriptions in order to preserve emotional signals. Notably, for image-based posts, we applied image-to-text conversion supplemented by human interpretation and prompt-based reasoning. These steps ensured that the final textual corpus was if excellent quality, semantically intact, and suitable for subsequent topic modeling and analysis.
@@ 3.3.1 Post-level Modeling
On the post level, given the limited number of posts (373) but their relatively longer length, we first used Latent Dirichlet Allocation (LDA) model from the Gensim library to extract latent topics. We tuned number_topics in the range of 5 to 30. The optimal number was 29 based on coherence scores (maximum of 0.44) and interpretability. The coherence value for a single topic measures how the top scoring words in this topic are semantically similar to each other, with higher coherence indicating that the topic is more interpretable and meaningful for human understanding Zhou et al. (2023).
@@
To better capture emotional sensitivity and health-related concerns beyond traditional narrative types, we also introduced two supplementary binary labels: grief signals and potential mental health risks. These were annotated based on explicit expressions of emotional loss, mourning, psychological distress, or self-identified vulnerability Cunsolo and Ellis (2018).

### v3

@@ 1 Introduction
Wildfires have intensified in intensity, severity, and duration on a global scale, largely due to the effects of climate change and increased human activity swain2025hydroclimate; mansoor2022elevation; yue2021assessing. These unplanned fire events result in widespread structural damage, economic losses, environmental degradation, and serious public health threats to affected communities bowman2017human; d2022wildfire; rappold2017community; wang2016spatial. The health risks associated with wildfires extend far beyond the immediate fire zones, encompassing not only physical and mental health concerns but also disparities in access to healthcare resources. Epidemiological and toxicological evidence links exposure to wildfire-derived fine particulate matter (PM2.5) to a range of symptoms, including respiratory and ocular irritation, and to adverse health outcomes such as asthma, cardiovascular diseases, negative birth outcomes, and skin inflammation zhang2025respiratory; lei2024wildfire. In addition, psychological impacts—such as anxiety, depression, and trauma-related disorders can result from smoke exposure, fear of loss, and evacuation-related stress lei2024wildfire; wettstein2024psychotropic. These mental health effects may persist long after fires are contained and can exacerbate physical health issues, particularly among vulnerable populations including children, older adults, pregnant individuals, those with preexisting cardiopulmonary conditions, and individuals from lower socioeconomic backgrounds lei2024wildfire; rappold2017community.
@@
When wildfires occur during irregular seasons and communities are unprepared, the consequences can be even more severe. In California’s Mediterranean climatic region, January has historically been a low-probability period for large-scale wildfires dong2022season. However, a hydroclimate whiplash homann2022linked; swain2025hydroclimate, driven by low humidity and hurricane-force Santa Ana winds, has induced abrupt changes from protracted wet periods to extreme dryness. This shift triggered the most devastating winter wildfires in the history. As of January 31, 2025, a total of 28 wildfires had burned 57,636 acres, resulted in 29 fatalities (pending coroner confirmation), destroyed 16,244 structures calfire_incidents, and forced the evacuation of approximately 200,000 residents StellohEtAl2025. Preliminary data indicate that the Eaton and Palisades fires in Los Angeles rank second and third in structural damage, and fifth and ninth in fatalities, respectively, though these figures remain provisional calfire_statistics. Despite the Storm Prediction Center SPC_FW2025, National Interagency Fire Center NIFC_outlook, and the local National Weather Service NWS_RFWLOX2025 had accurately forecasted the windstorm and heightened fire risk, the actual spread and scale of wildfires far exceeded expectations.
@@
Amid rapidly evolving wildfire conditions, many individuals have turned to social media platforms—such as Twitter and Reddit—for hyperlocal information, health-related resources, and mutual aid slavkovikj2014review; lever2022social; ma2024investigating. While Twitter has been extensively studied in the context of disaster response, Reddit’s role in collective information sharing and community support remains comparatively underexplored chen2019online. Traditional data sources, including clinical records and post-disaster surveys, offer valuable medical insights into wildfire-related health outcomes; however, they often suffer from limitations such as small sample sizes, limited scalability, high costs, and delayed data availability garcia2024wildfires. With the California fire season approaching, there is an urgent need for rapid, scalable methods to detect public health signals and uncover unmet needs in affected communities in real time.
@@
Social media platforms serve as critical sources of information during disasters, significantly supporting disaster response through the rapid dissemination of real-time updates ye2020social. Given the sheer volume of data generated on social media platforms far beyond the capabilities of manual examination, existing research leverages machine learning techniques, such as keyword-based methods or classification models, which utilize textual and metadata features, to identify and categorize relevant social media content vongkusolkit2021situational; huang2015geographic. Most recently, pretrained large language models (LLMs) have become the state-of-the-art method in natural language processing (NLP). Built on top of these pretrained embeddings, supervised and unsupervised methods are applied as post-processing strategies. Supervised techniques still require a certain quantity of annotated data for fine-tuning, which can be time-consuming in fast-moving emergencies. By contrast, unsupervised methods, such as topic modeling, can quickly identify latent topics based on semantic clustering without annotation efforts zhou2023guided. Furthermore, most previous research on disaster events has considerable time lags, with findings published years after the crisis. These findings can inform long-term mitigation, but often lack value for immediate response.
@@
In order to quickly identify and extract actionable information for disaster response and management, a topic classification schema or a coding schema is needed to guide the categorization of social media datasets vongkusolkit2021situational; verma2011natural; hughes2009twitter. These guidance schema often include categories such as casualties, damage reports, donation requests, or public emotions. However, designing an effective classification schema and corresponding information extraction methods is complex and highly context-dependent, influenced by factors such as (1) the type of disaster event, (2) the specific goals of the study or analysis, and (3) the characteristics of the social media platform used vongkusolkit2021situational.
@@ 2 Literature Review
Topic modeling is an unsupervised NLP technique that enables the identification and organization of latent thematic structures within large text corpora. This approach is particularly valuable for managing the exponential growth of digital textual data—from emails and reports to social media posts—by converting unstructured content into structured, analyzable representations vayansky2020review.
@@ 2.1 Topic Modeling Algorithms
The core of many topic modeling methods is the “bag of words” paradigm, where the focus is on word frequencies rather than their order zhang2010understanding. This paradigm generates two critical matrices: a word-topic matrix that encapsulates topics through their associated words, and a document-topic matrix that reflects each document’s composition in terms of these topics. Such representations help extract important insights from vast corpora, especially in the absence of explicit labeling barde2017overview.
Historically, topic modeling has evolved significantly—from early models such as Latent Semantic Analysis (LSA) deerwester1990indexing and its probabilistic extension, Probabilistic LSA (pLSA) hofmann1999probabilistic, to more sophisticated Bayesian approaches like the Gibbs sampling algorithm for the Dirichlet multinomial mixture (GSDMM) yin2014dirichlet and Latent Dirichlet Allocation (LDA). LDA, for instance, leverages Dirichlet distributions to iteratively refine the assignment of words to topics based on the probability of topic occurrence in documents and word occurrence within topics blei2003latent. Concurrently, Non-Negative Matrix Factorization (NMF) emerged as an alternative linear algebra-based method that decomposes high-dimensional text data into interpretable, lower-dimensional representations lee1999learning.
In recent years, the advent of transformer-based models has ushered in a new era of topic modeling with algorithms such as Top2Vec angelov2020top2vec and BERTopic grootendorst2022bertopic. Top2Vec detects topics with assumption that each document is primarily centered on a single topic. In contrast, BERTopic performs powerful and flexible modular structure by integration of sophisticated sentence embedding techniques, dimensionality reduction, and clustering algorithms for modern text analytics.
Overall, the evolution from LSA and pLSA to GSDMM, LDA, NMF, and the transformer-based methods illustrates the dynamic development of topic modeling methodologies. Each algorithm has its own merits for applications depending on the trade-offs among scalability, interpretability, and contextual sensitivity. For example, LDA is well-suited for longer texts but is typically efficient only when applied to a relatively small number of documents. In contrast, BERTopic performs better on shorter texts but requires a large corpus to effectively capture and cluster meaningful topics egger2022topic.
@@ 2.2.1 Dataset Sources
Structured data sources, such as NASA’s Aviation Safety Reporting System (ASRS) and the ICS-209-PLUS incident database, have been used in a limited number of studies focusing on operational safety and technical failures in wildfire contexts mbaye2023bert; andrade2023machine). While these datasets are valuable in terms of infrastructure performance and risk reporting, they are less suited for examining public sentiment, communication, and response in real time.
In contrast, most studies rely on social media data, particularly from Twitter, to analyze wildfire trends, public discourse, and crisis communication strategies. These datasets vary in size, ranging from thousands to millions of posts, and span from short-term event monitoring to multi-year longitudinal analyses zander2023trends; garcia2024wildfires; souza2024using; madichetty2021multi; madichetty2023roberta. For instance, García et al. (2024) focused on mental health during the 2017 Tubbs Fire in California, while Zander et al. (2023) examined broader situational awareness during Australia’s “Black Summer” bushfires in 2019 and 2020.
Due to limitations associated with Twitter data—such as character constraints, low information density, credibility issues, and recent restrictions on data access—researchers have begun to explore alternative platforms. Reddit, with its longer-form user content, community-based moderation, and more accessible data policies, presents a compelling alternative for analyzing public discourse around wildfires lever2022social.
@@ 2.2.2 Topic Classification Schema
After latent topic clusters are generated, it is common to organize them into interpretable categories using domain annotations. Many studies implicitly follow two analytical rationales: situational awareness (SA) and emotional response. SA is defined as “all knowledge that is accessible and can be integrated into a coherent picture, when required, to assess and cope with a situation” sarter2017situation, derived from three-level model of perception, comprehension, and projection endsley1995measurement. In the context of disaster response, SA captures real-time information such as fire progression, evacuation orders, and resource availability. While SA-based categories are prevalent across studies, shown in Table 1, they tend to vary widely ranging from three to fifteen categories depending on research goals. This inconsistency reflects a lack of unified structure for organizing real-time information in disaster response.
To capture emotional response, many studies rely on sentiment analysis lever2022social; garcia2024wildfires; zander2023trends, which classifies corpus categorizing text into positive, negative, or neutral categories, as well as fundamental emotion such as fear, anger, or sadness. While sentiment analysis provides general emotional trends, it is often insufficient to summarize intricate expressions such as narratives of blame, resilience, or justice ruiz2017twitter. Moreover, sentiment labels may be inaccurate in informal, sarcastic, or ambiguous media messages wankhade2022survey.
@@
madichetty2021multi Califonia wildfires Twitter 2 weeks 15 Fine-tuned BERT
lever2022social California Wildfire Season (2016) Twitter, Reddit 4 months 5 LDA
ma2024investigating Western Wildfire Season(2020) Twitter 1 month 6 BERTopic
zander2023trends Black Summer(2019-2020) Twitter 6 months 8 GSDMM
souza2024using Wildfires in Brazil National Park Twitter 11 years 6 BERTopic
garcia2024wildfires Tubbs fire Twitter 3 weeks 8 LDA
andrade2023machine Wildfire incidents ICS-209-PLUS 8 years 3 BERTopic
mbaye2023bert Wildfire reports UAS 4 years 7 BERTopic
To address these limitations, we draw on the concept of Crisis Narratives from crisis communication research. Crisis Narratives (CN) refer to “the stories told about crises that carry meaning, encode lessons, and frame larger public and societal understanding of risks, warnings, and potential harm” seeger2016narratives; liu2020telling. Common narrative themes, such as blame, victimhood, heroism, and renewal, have been widely used to examine institutional rhetoric and public interpretation in disaster contexts, as shown in table 2. Despite their relevance, this framework has not yet been systematically applied to social media data in the context of wildfires. In this study, we adopt and extend CN theory to move beyond surface-level sentiment analysis, offering a more nuanced and structured lens for interpreting emotionally charged public discourse during crises.
@@
wolfe2016organizing Disaster Interview, participant observation, public meetings, media reports and organizational documents Blame and Accountability, Victim-Centered Disaster, Renewal and Growth
liu2020telling Pandemic Quantitative survey Blame, Renewal, Victim, Hero, Memorial
an2024dynamics Pandemic Twitter Blame, Renewal, Victim, Hero, Memorial
@@ 3.2 Data Preprocessing
We adapted a multi-step preprocessing to clean the Reddit data for analysis. In the first step, we removed special characters, HTML tags, and Markdown symbols. We excluded instances with fewer than 10 words based on the recommended effective semantic length (15-20 words) markel2009technical. Unlike traditional NLP methods that completely delete emojis and URLs, we extracted and saved 1,002 unique URLs embedded in posts and comments for rank analysis. Emojis were converted to their corresponding textual descriptions in order to preserve emotional signals. Notably, for image-based posts, we applied image-to-text conversion supplemented by human interpretation and prompt-based reasoning. These steps ensured that the final textual corpus was if excellent quality, semantically intact, and suitable for subsequent topic modeling and analysis.
@@ 3.3.1 Post-level Modeling
On the post level, given the limited number of posts (373) but their relatively longer length, we first used Latent Dirichlet Allocation (LDA) model from the Gensim library to extract latent topics. We tuned number_topics in the range of 5 to 30. The optimal number was 9 based on coherence scores (maximum of 0.49) and interpretability. The coherence value for a single topic measures how the top scoring words in this topic are semantically similar to each other, with higher coherence indicating that the topic is more interpretable and meaningful for human understanding zhou2023guided.
@@
To better capture emotional sensitivity and health-related concerns beyond traditional narrative types, we also introduced two supplementary binary labels: grief signals and potential mental health risks. These were annotated based on explicit expressions of emotional loss, mourning, psychological distress, or self-identified vulnerability cunsolo2018ecological.